OpenStudy (anonymous):

calculate the indefinite integral of ∫(ln(acos(x)) dx)/(acos(x) √(1-x^2 ))

OpenStudy (anonymous):

all yours!

OpenStudy (anonymous):

its a setup!

OpenStudy (anonymous):

oh so it is

OpenStudy (anonymous):

I got it on a homework ;(

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\ln(a \cos(x) ) dx}{a \cos(x) \sqrt{1-x^2}}\]

OpenStudy (anonymous):

im assuming thats arccos.

myininaya (myininaya):

so i have the problem right problem?

OpenStudy (anonymous):

i think it is \[arc\cos(x)\] rather than \[acos(x)\] makes it trivialish

myininaya (myininaya):

oh i thought a was a constant

OpenStudy (anonymous):

i did too, but that sqrt at the end sorta screams "setup!!"

OpenStudy (anonymous):

yea lol, srry, i ushually use to the negative 1st power

myininaya (myininaya):

if so try letting u=arccos(x)

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

i was hoping i missed something. or is my math teacher just being mean?

OpenStudy (anonymous):

let u = ln(arccosx), then du would be 1/arccosx . 1/sqrt(1-x^2)

OpenStudy (anonymous):

right? or am i getting my derivative of arccos wrong again >.>

OpenStudy (anonymous):

because i hate calculus <.<

OpenStudy (anonymous):

no it is right and myininaya is about to prove it again i can tell

OpenStudy (anonymous):

i think its negative

myininaya (myininaya):

\[\int\limits_{}^{} \frac{\ln(u)}{u} (-du)\] let v=ln(u) then dv=1/u du \[-\int\limits_{}^{} v dv\]

myininaya (myininaya):

is it really that easy or did i make mistake?

OpenStudy (anonymous):

yes is it a setup as joemath said

OpenStudy (anonymous):

its a TRAP!

OpenStudy (anonymous):

tarp it's a tarp

myininaya (myininaya):

\[-\frac{v^2}{2}+C=-\frac{[\ln(u)]^2}{2}+C=-\frac{[\arccos(x)]^2}{2}+C\]

OpenStudy (anonymous):

now you messed up

myininaya (myininaya):

oops i forgot the natural log

OpenStudy (anonymous):

forgot the log!

myininaya (myininaya):

\[\frac{-[\ln|\arccos(x)|]^2}{2}+C\]

myininaya (myininaya):

how satellite i thought i was going to prove something i'm unpredictable haha

OpenStudy (anonymous):

bam.

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

lol thanks for helping me maintain my 100% :)

OpenStudy (anonymous):

damn. i guess this time it was ok to remember...

myininaya (myininaya):

lol

myininaya (myininaya):

this is the part where i told to class i showed this many times so you do it on your own now have fun

OpenStudy (anonymous):

averaging less than 4 hours of sleep really isn't good for doing math. for which i blame the IB diploma program

OpenStudy (anonymous):

hate calculus >.> hey satellite, for that problem i brought up last night, apparently we were supposed to guess what form the anti-derivative was in, and solve a system of equations to show it was rational >.>

myininaya (myininaya):

omg i never would have thought that joe for your problem i looked at too

OpenStudy (anonymous):

i dont know how anyone is supposed to guess that, but whatev >.<

myininaya (myininaya):

all of my sentences are coming out incomplete its like i don't read before i click post

OpenStudy (anonymous):

what? what system of equations?

myininaya (myininaya):

zarkon deleted is post lol

myininaya (myininaya):

his*

OpenStudy (anonymous):

like, it was basically:\[\int\limits\frac{x^4-2x^3+3x^2-2x+1}{(x^3-3x+1)^2}\]and we are supposed to say "that kinda looks like a quotient rule with a 2nd degree polynomial over a 3 degree polynomial" >.>

myininaya (myininaya):

i finally win

OpenStudy (anonymous):

what did you win? a cookie?

myininaya (myininaya):

yes give it back please

OpenStudy (anonymous):

well it does look like a quotient rule problem because the denominator is something squared

OpenStudy (anonymous):

so we make a guess that:\[F(x)=\frac{ax^2+bx+c}{x^3-3x+1}\]take the derivative, and solve for a b and c. it gives a = -1, b = 1, and c = 0.

myininaya (myininaya):

thats kindof neat that derivative is nasty though

OpenStudy (anonymous):

yeah =/ but it worked lol. now i have to work on making my solution look neater and more understandable.

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

hate calculus >.>

OpenStudy (anonymous):

yeah it pretty much s*cks i agree.

myininaya (myininaya):

i can't say differentiate very well so i say take derivative

myininaya (myininaya):

why they make such big words i'm almost blonde come on

OpenStudy (anonymous):

almost blond? lol and the worst thing about calculus is it is endless

myininaya (myininaya):

i use to not be able to say statistics without stuttering

OpenStudy (anonymous):

calc 1 calc 2 calc 3 calc 10 calc theta etc etc

OpenStudy (anonymous):

dont forget pre-calc >.> lolol

OpenStudy (anonymous):

i bet you can't say 'stuttering statistics" without stuttering

myininaya (myininaya):

lol